We showed that if S and T are isomorphic groups and e is the identity for S ,φ(e) gives the identity for T. Now show that if a and a′ are inverses in S, then φ (a) and φ (a′) are inverses in T


. (In other words, show that φ (a′) = (φ(a))'


1
Expert's answer
2021-02-24T12:41:45-0500

Solution:

We have, a.a'=e. (since a and a' are inverse)

=> φ (a.a′) =φ (e) (applying φ to both sides)

=> φ (a).φ (a′) =φ (e) (as φ is isomorphism)

=> φ (a).φ (a′) =φ (a).[φ (a)]' (as φ(e) is identity of T)

=> φ (a′) =[φ (a)]' (left cancelling φ (a))

Hence proved.